OpenStudy (anonymous):

SEQUENCE PROBLEMS WILL FAN AND MEDAL these are all for a study guide 1. find the next four terms of the sequence below: 24, 19.2, 15.36, _____, _____, _____, _____ what is the rude for this sequence? 2. what is the 17th term of the sequence that begins -3, 9, -27, 81...? 3. the fifth term of a geometric sequence is 120, and the seventh term is 30. What are the possible values of the sixth term of the sequence?

OpenStudy (anonymous):

do you go to connections 8grade

OpenStudy (anonymous):

here's the questions

OpenStudy (anonymous):

@moklsss im a senior at connections this is the algebra 2 final study guide

OpenStudy (anonymous):

do u know anyone 8grade

OpenStudy (anonymous):

@johnweldon1993 these are the other ones i don't remember how to find the rule, i know there was a trick aside from basically counting it out but i can't remember

OpenStudy (johnweldon1993):

These are arithmetic? I think these are geometric arent they?

OpenStudy (anonymous):

quite possibly honesly

OpenStudy (anonymous):

i just know i have to find the n= rule for the sequences and then you plug in the nth number that you need to find but i dont remember how to find the rule

OpenStudy (johnweldon1993):

Lol I believe they are :D Okay so fir the first one... 24, 19.2, 15.36, ____ , ____ , ____ , ____ To find out *because this is geometric* the next terms...we divide the second number, by the first number....or the third number by the 2nd number...either way you will get the same "common ratio" Then yes you can just plug it into \[\large x_n = ar^{x - 1}\]

OpenStudy (anonymous):

check ur msgggg @jesselaurenx3 can ur sis help me"?

OpenStudy (anonymous):

@moklsss she is done with school and totally on vacation i doubt it

OpenStudy (anonymous):

okay so xn=arx−1 would be the rude? and then so how would i find the next four with that?

OpenStudy (anonymous):

@jesselaurenx3 do u have her lgogin to help me?

OpenStudy (anonymous):

OH OKAY I FIGURED THIS OUT theres a .8 difference so the next 4 would be 9.83, 7.86, 6.29, 5.03 because when i divided 19.9/24 i got 0.8 and so on so i multiplied the next number by .8 to get the next four not sure what that makes the rule?

OpenStudy (johnweldon1993):

Okay so with the equation \[\large x_n = ar^{n-1}\] a = the first term of the sequence r = the common ratio you still need to find above :P And 'n is the term you are solving for

OpenStudy (johnweldon1993):

Beat me to it >.< :P lol but as far as those results...hmm Well for the 4th term...I get \[\large x_4 = 24(0.8)^{4 - 1}\] which comes out to

OpenStudy (anonymous):

yeah, not sure if my calculator came out to the right things

OpenStudy (johnweldon1993):

So what I get is \[\large x_4 = 24(0.8)^{4-1} = 12.288\] \[\large x_5 = 24(0.8)^{5-1} = 9.8304\] \[\large x_6 = 24(0.8)^{6-1} = 7.86432\] \[\large x_7 = 24(0.8)^{7-1} = 6.291456\]

OpenStudy (anonymous):

right, i just got that backwards somehow

OpenStudy (johnweldon1993):

Not a problem :) #'s 8 and 9 on your worksheet are the same process as well...sadly I must go now :( But you got this :D if you need anything else just message me and I'll get back to you as soon as I can :)

OpenStudy (anonymous):

alrighty! thanks!